### v1

@@ Abstract
3D human pose estimation, knowledge distillation.
@@ I Introduction
3D Human Pose Estimation (HPE) detects 3D joint coordinates from monocular video and is widely used in downstream tasks like action recognition [1, 2] and motion analysis [3, 4]. To achieve high accuracy, prior methods typically use multi-frame input sequences to capture temporal dependencies [5, 6, 7, 8, 9, 10, 11, 12] and large networks (e.g., Vision Transformers [13]) for feature extraction. However, these approaches are computationally expensive and suffer from slow inference speeds due to dense inputs and significant network overhead.
@@
Figure 1: Comparison between the previous methods PoseFormerV2 [15], HDFormer [8], MHFormer [10] and our SCJD on Human3.6M. $f$ represents the length of the input sequence. For both metrics, the smaller Mean Per Joint Position Error (MPJPE) and floating-point operations (FLOPs), the better.
To address these challenges, we propose a novel Sparse Correlation and Joint Distillation (SCJD) for 3D HPE. Our key insight is to leverage sparse input sequences and KD to reduce computational overhead while maintaining accuracy. Specifically, Our approach begins with Sparse Correlation Input Sequence Downsampling, which sparsifies the student network’s inputs to reduce redundancy while preserving inter-frame correlations more effectively than single-frame methods. Next, to transfer knowledge effectively from a dense, heavy teacher network to a sparse, lightweight student network, we introduce Dynamic Joint Spatial Attention Distillation, which comprises two components: a) Dynamic Joint Embedding Distillation, which aggregates the local context information of teacher’s spatial encoder to enhance the student’s spatial feature representation. b) Adjacent Joint Attention Distillation, which employs an adjacent joint attention matrix to guide the student network in focusing on connections between adjacent joints, improving its understanding of human posture. Finally, we propose Temporal Consistency Distillation, which aligns the temporal correlation between input contexts by upsampling the student’s features to match the teacher network’s outputs, and complemented by global supervised learning with ground truth. Our extensive experiments validate our method’s superiority over existing state-of-the-art techniques in terms of both efficiency and accuracy.
To summarize, our contributions are threefold:
@@
Figure 2: Overview of the SCJD framework: SCJD comprises two transformer encoders that separately learn spatial and temporal information. The student model first employs Sparse Correlation Input Sequence Downsampling to reduce input redundancy while preserving temporal correlations. Then, through Dynamic Joint Embedding Distillation and Adjacent Joint Attention Distillation, knowledge from the teacher network’s spatial encoder is transferred to help the student model learn joint associations. Finally, Temporal Consistency Distillation facilitates the transfer of temporal knowledge from the teacher’s temporal encoder, enabling the student model to capture the temporal relationships between multi-frame postures.
@@ II-B Knowledge Distillation in 3D Human Pose Estimation
Knowledge distillation [17] is an effective technique for compressing networks by transferring knowledge from a large teacher network to a lightweight student network using deep learning technology [22, 23], improving the latter’s accuracy. Several studies [18, 19, 20, 24, 25] have applied distillation to accelerate 3D human pose estimation. Hwang et al. [18] trained a student network to mimic the teacher’s output logits, while Kim et al. [19] used pruning and feature imitation to boost performance. Bian et al. [20] transferred multi-view stereo knowledge to the student model. However, these methods are limited to single-frame inputs and fail to utilize multi-frame video context, which is crucial for resolving 2D-to-3D ambiguities. Additionally, they overlook the importance of teaching the student network to learn joint spatial relationships. Our approach addresses these gaps by designing a distillation framework that incorporates multi-frame context while emphasizing joint relationships.
@@ III-C Dynamic Joint Spatial Attention Distillation
Adjacent Joint Attention Distillation. After aligning the joint embedding feature between the teacher and student networks, we initiate knowledge transfer from the teacher to the student. However, simply replicating the teacher network’s feature extraction capabilities makes it difficult for the student network to grasp the underlying relationships between human joints. Note that human posture is inherently represented as a skeletal model, where interconnected joints are linked by bones, forming a robust prior topological structure, as shown in Fig. 2(d). Adjacent joints are defined as pairs of keypoints directly connected within this structure, thereby capturing the local spatial relationships of the human pose. The interactions between adjacent joints not only encode essential structural characteristics of the skeleton but also provide valuable spatial-temporal constraints and biomechanical priors for the model. These connectivities between adjacent joints are crucial prior knowledge that guides the student network to focus on these relationships, thereby enhancing its fundamental understanding of human posture. To incorporate this structural prior, we design an adjacent joint mask $\mathcal{M}_{mask}$, which enforces the student network to learn the teacher network’s multi-head self-attention (MHSA) patterns specifically between adjacent joints. This $\mathcal{M}_{mask}$ can be expressed as:
$\mathcal{M}_{mask}(i,j)=\begin{cases}1,&\text{if bone }i\text{ is connected tobone }j,\\ 0,&\text{otherwise}.\end{cases}$ (2)
To encourage the student network to focus on adjacent joint features derived from the teacher network, we multiply $\mathcal{M}_{mask}$ with the MHSA matrix generated by both teacher and student network’s spatial encoder. This results in the adjacent joint attention $\mathcal{M}$, which filters out irrelevant elements in the attention matrix. The computation is expressed as:
$\mathcal{M}(\mathcal{F})=\frac{\mathcal{F}\mathcal{F}^{tr}}{\sqrt{C}}\otimes\mathcal{M}_{mask},$ (3)
@@ IV-A Experimental Settings
Implementation Details. The input sequence lengths of the student network are $f^{S}=\left\{3,9,27\right\}$. For the teacher network, we chose $f^{T}=81$. The embedding dimensions of the student and teacher network are $C^{S}=16$ and $C^{T}=32$. The number of joints $J=17$. In dynamic token matching, we set $N=20$, $K=3$. We empirically set the hyperparameters of Eq. 8 to $\alpha=\beta=1$, $\gamma=0.01$. We trained our model for 50 epochs using the Adam optimizer with learning rate of 1e-3 and weight decay of 0.1. For the network structure and other more details, please refer to the supplementary material.
@@ IV-C Comparisons with Knowledge Distillation Methods
TABLE II: Comparisons with Previous Low-frame-rate Methods on MPI-INF-3DHP. Using ground-truth 2D poses as inputs. Best in red, second best in blue.
@@
TABLE III: Comparisons with Traditional KD Methods on Human3.6M. The 2D poses obtained by CPN [29] are used as inputs. Best in red, second best in blue.
@@
TABLE IV: Ablation Study on Different Parts on Human3.6M
@@
TABLE V: Ablation Study on Distillation Components on Human3.6M
@@
TABLE VI: Ablation Study on Parameter Settings on Human3.6M
@@
IV-D Ablation Study
Hyperparameter Analysis. We have further analyzed the hyperparameters of the loss function (Eq. 8), with the corresponding results summarized in Table VII. Specifically, we examine the weights $\alpha$, $\beta$, and $\gamma$, which correspond to the individual loss components. The results indicate that our method achieves the best performance when $\alpha=\beta=1$ and $\gamma=0.01$. We attribute this to the fact that these settings effectively scale the losses to the same order of magnitude, thereby facilitating more stable and balanced optimization. We will add this analysis and discussion in the revision.
V Conclusion
Supplementary Material
The supplementary material contains: 1) Additional Details of the Network Structure; 2) More Experimental Details; 3) Qualitative Comparisons on In-the-wild Videos; 4) More Qualitative Comparisons on Human3.6M.
Figure 4: Left: details of the spatial transformer. Center: details of the temporal transformer. Right: details of the transformer encoder structure.
VI Additional Details of the Network Structure
In our framework, spatial transformers, as shown in Fig. 4(a), are used to model the positional connections among joints in a frame of pose. Building on PoseFormer [26], for each frame of the input 2D pose sequence, coordinates of all joints of a pose are linearly projected to a higher-dimensional space, and a learnable spatial positional embedding is added to encode the positions of the joints. These features are then passed through a spatial transformer encoder. Then, the output of each frame of the spatial transformer encoder is concatenated as the input of the temporal transformer encoder, as shown in Fig. 4(b). Temporal transformer encoders capture frame-to-frame dependencies across the sequence, and its output is passed through the regression head to obtain the 3D pose for the center frame.
The spatial transformer encoder and temporal transformer encoder in our framework both adopt the standard transformer encoder architecture [39], as illustrated in Fig. 4(c). This architecture consists of N identical layers, each comprising two main components: a Multi-Head Self-Attention (MHSA) mechanism and a Feed Forward Network (FFN). Layer Normalization (Layer Norm) is applied before the MHSA and FFN to stabilize training, while residual connections are employed around these sublayers to improve gradient flow and facilitate efficient learning. The MHSA module enables the encoder to capture complex dependencies within the input sequences, while the FFN refines the intermediate representations.
Figure 5: Qualitative comparisons between ours and PoseFormerV2 [15] on in-the-wild videos. The same visualization setting as PoseFormerV2 [15] is adopted, the 2D pose is obtained by the detector HRNet [40], and the number of input frames of the model is f=27. The black arrows highlight locations where our method clearly has better results.
VII More Experimental Details
Implementation Details. The input sequence length of the teacher model is $f_{T}=81$. For the student model, we chose three different frame sequence lengths, i.e., $f_{S}=3$, $f_{S}=9$, and $f_{S}=27$. Pose level flipping was applied as a data augmentation in training and testing [41]. We trained our model using the Adam [42] optimizer for 50 epochs with a weight decay of 0.1. We adopt an exponential learning rate decay schedule with an initial learning rate of 1e-3 and a decay factor of 0.98 for each epoch. An NVIDIA GeForce RTX 4090 GPU is used for training with the batch size set to 1024. The details about model parameters of teacher network and student network with results are discussed in the ablation studies (Sec. IV-C of the article: The Smaller Model).
VIII Qualitative Comparisons on In-the-wild Videos
IX More Qualitative Comparisons on Human3.6M

### v2

@@ Abstract
3D Human Pose Estimation, distillation.
@@ I Introduction
3D Human Pose Estimation (HPE) detects 3D joint coordinates from monocular video and is widely used in downstream tasks like action recognition [1, 2] and motion analysis [3, 4]. To achieve high accuracy, prior methods typically use multi-frame input sequences to capture temporal dependencies [5, 6, 7, 8, 9, 10, 11] and large networks (e.g., Vision Transformers [12]) for feature extraction. However, these approaches are computationally expensive and suffer from slow inference speeds due to dense inputs and significant network overhead.
@@
Figure 1: Comparison between the previous methods PoseFormerV2 [14], HDFormer [8], MHFormer [10] and our SCJD on Human3.6M. $f$ represents the length of the input sequence.
To address these challenges, we propose a novel Sparse Correlation and Joint Distillation (SCJD) for 3D HPE. Our key insight is to leverage sparse input sequences and KD to reduce computational overhead while maintaining accuracy. Specifically, Our approach begins with Sparse Correlation Input Sequence Downsampling, which sparsifies the student network’s inputs to reduce redundancy while preserving inter-frame correlations more effectively than single-frame methods. Next, to transfer knowledge effectively from a dense, heavy teacher network to a sparse, lightweight student network, we introduce Dynamic Joint Spatial Attention Distillation, which comprises two components: a) Dynamic Joint Embedding Distillation, which aggregates the local context information of teacher’s spatial encoder to enhance the student’s spatial feature representation. b) Adjacent Joint Attention Distillation, which employs an adjacent joint attention matrix to guide the student network in focusing on connections between adjacent joints, improving its understanding of human posture. Finally, we propose Temporal Consistency Distillation, which aligns the temporal correlation between input contexts by upsampling the student’s features to match the teacher network’s outputs, and complemented by global supervised learning with ground truth. Our extensive experiments validate our method’s superiority over existing state-of-the-art techniques in terms of both efficiency and accuracy. Our contributions are threefold:
@@
Figure 2: Overview of the SCJD framework. SCJD comprises two transformer encoders that separately learn spatial and temporal information. The student model first employs Sparse Correlation Input Sequence Downsampling to reduce input redundancy while preserving temporal correlations. Then, through Dynamic Joint Embedding Distillation and Adjacent Joint Attention Distillation, knowledge from the teacher network’s spatial encoder is transferred to help the student model learn joint associations. Finally, Temporal Consistency Distillation facilitates the transfer of temporal knowledge from the teacher’s temporal encoder, enabling the student model to capture the temporal relationships between multi-frame postures.
@@ II-B Knowledge Distillation in 3D Human Pose Estimation
Knowledge distillation [16] is an effective technique for compressing networks by transferring knowledge from a large teacher network to a lightweight student network, improving the latter’s accuracy. Several studies [17, 18, 19] have applied distillation to accelerate 3D human pose estimation. Hwang et al. [17] trained a student network to mimic the teacher’s output logits, while Kim et al. [18] used pruning and feature imitation to boost performance. Bian et al. [19] transferred multi-view stereo knowledge to the student model. However, these methods are limited to single-frame inputs and fail to utilize multi-frame video context, which is crucial for resolving 2D-to-3D ambiguities. Additionally, they overlook the importance of teaching the student network to learn joint spatial relationships. Our approach addresses these gaps by designing a distillation framework that incorporates multi-frame context while emphasizing joint relationships.
@@ III-C Dynamic Joint Spatial Attention Distillation
Adjacent Joint Attention Distillation. After aligning the joint embedding features of the teacher and student networks, we initiate knowledge transfer. However, merely replicating the teacher network’s feature extraction process is insufficient for the student network to grasp the underlying relationships between human joints. Human posture is naturally represented as a skeletal model, where interconnected joints form a structured topology, as illustrated in Fig. 2(d). In this structure, adjacent joints are directly connected keypoints, capturing local spatial relationships crucial to human pose representation. These interactions encode essential skeletal characteristics and provide valuable spatial-temporal constraints and biomechanical priors for the model. The connectivity of adjacent joints serves as critical prior knowledge, guiding the student network to focus on these relationships and enhancing its understanding of human posture. To incorporate this structural prior, we introduce an adjacent joint mask $\mathcal{M}_{mask}$, which directs the student network to learn the teacher network’s patterns between adjacent joints. This $\mathcal{M}_{mask}$ is defined as:
$\mathcal{M}_{mask}(i,j)=\begin{cases}1,&\text{if bone }i\text{ is connected tobone }j,\\ 0,&\text{otherwise}.\end{cases}\vspace{-8pt}$ (2)
To encourage the student network to focus on adjacent joint features derived from the teacher network, we multiply $\mathcal{M}_{mask}$ with the attention matrix generated by both teacher and student network’s spatial encoder. This results in the adjacent joint attention $\mathcal{M}$, which filters out irrelevant elements in the attention matrix. The computation is expressed as:
$\mathcal{M}(\mathcal{F})=\frac{\mathcal{F}\mathcal{F}^{tr}}{\sqrt{C}}\otimes\mathcal{M}_{mask},\vspace{-5pt}$ (3)
@@ IV-A Experimental Settings
Implementation Details. The input sequence lengths of the student network are $f^{S}=\left\{3,9,27\right\}$. For the teacher network, we chose $f^{T}=81$. The embedding dimensions of the student and teacher network are $C^{S}=16$ and $C^{T}=32$. The number of joints $J=17$. In dynamic token matching, we set $N=20$, $K=3$. We empirically set the hyperparameters of Eq. 8 to $\alpha=\beta=1$, $\gamma=0.01$. We trained our model for 50 epochs using the Adam optimizer [24, 25, 26] with learning rate of 1e-3 and weight decay of 0.1. For the network structure and other more details, please refer to the supplementary material.
@@ IV-C Comparisons with Knowledge Distillation Methods
TABLE II: Comparisons with Previous Low-frame-rate Methods on MPI-INF-3DHP. Using ground-truth 2D poses as inputs.
@@
TABLE III: Comparisons with Traditional KD Methods on Human3.6M. The 2D poses obtained by CPN [27] are used as inputs.
@@
TABLE IV: Ablation Study on Different Parts on Human3.6M.
@@
TABLE V: Ablation Study on Distillation Components on Human3.6M.
@@
TABLE VI: Ablation Study on Parameter Settings on Human3.6M.
@@
IV-D Ablation Study
V Conclusion
Appendix A More Experimental Details
Implementation Details. The input sequence length of the teacher model is $f_{T}=81$. For the student model, we chose three different frame sequence lengths, i.e., $f_{S}=3$, $f_{S}=9$, and $f_{S}=27$. Pose level flipping was applied as a data augmentation in training and testing [39]. We trained our model using the Adam optimizer for 50 epochs with a weight decay of 0.1. We adopt an exponential learning rate decay schedule with an initial learning rate of 1e-3 and a decay factor of 0.98 for each epoch. An NVIDIA GeForce RTX 4090 GPU is used for training with the batch size set to 1024. The details about model parameters of teacher network and student network with results are discussed in the ablation studies (Sec. IV-C of the article: The Smaller Model).
A-A Qualitative Comparisons on In-the-wild Videos
A-B More Qualitative Comparisons on Human3.6M
A-C Hyperparameter Analysis
We have further analyzed the hyperparameters of the loss function (Eq. 8), with the corresponding results summarized in Table VII. Specifically, we examine the weights $\alpha$, $\beta$, and $\gamma$, which correspond to the individual loss components. The results indicate that our method achieves the best performance when $\alpha=\beta=1$ and $\gamma=0.01$. We attribute this to the fact that these settings effectively scale the losses to the same order of magnitude, thereby facilitating more stable and balanced optimization.